OpenStudy (anonymous):

4pencils and five pens costs $2.00 Three pencils and four pens cost $1.58 find the cost of a pencil and the cost of a pen

OpenStudy (amistre64):

L = lead P = pen 4L + 5P = 2 3L +4P = 1.58

OpenStudy (anonymous):

asaila1 use simultaneous equations

OpenStudy (amistre64):

L = 2/4 -5P/4

OpenStudy (anonymous):

you should get pen = $0.24 and pencil =$0.20

OpenStudy (amistre64):

we could multiply the top by 3 and the bottom by -4;)

OpenStudy (amistre64):

12L +15P = 6 -12L -16P = -6.32 --------------- 0 -P = -.32 P = .32 ???

OpenStudy (amistre64):

L = .10? lets see :) 3(.1) +4(.32) = 1.58 .3 + 1.28 = 1.58 it works

OpenStudy (amistre64):

pencil = .10 pen = .32

OpenStudy (amistre64):

vit; how you get yours?

OpenStudy (anonymous):

let p = pencil and x = pens so 4p + 5x =$2.00---- 1 and 3p +4x =$1.58---- 2 make p the subject in equation 1: p = (2-5x)/ 4 sub p into equation 2 3((2-5x)/ 4) + 4x =$1.58 ((6-15x)/4)+4x =$1.58 x+6 = $6.24 x= $0.24 sub x back into p p = (2-5($0.24))/4 p = $0.20

OpenStudy (anonymous):

ooops i am so sorry amistre64 you are correct i made a slight error in my calculations in my working out on paper i wrote $1.56 MY APOLOGIES I AM SO SORRY!

OpenStudy (amistre64):

virtus; its ok :) happens to me ALLLL the time ;) its always good to have another set of eyes :)

OpenStudy (anonymous):

=)